OpenStudy (anonymous):

In which expression should the exponents be multiplied? one fifth to the 2nd times one fifth to the 6th 9 to the 3rd over 9 to the 4th 73 ⋅ 78 (26)−5

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

after this there is 2 more

OpenStudy (phi):

you shout use ^ or the equation editor. I can't tell which numbers are exponents.

OpenStudy (anonymous):

oh let me fix it

OpenStudy (phi):

but we did this in your first question \[ \left(7^3\right)^2 = 7^6\] notice that 2*3= 6 (that is the short cut)

OpenStudy (anonymous):

1/5^2 x 1/5^6 9^3/9^4 7^3 x 7^8 (2^6)^-5

OpenStudy (anonymous):

do u want a screenshot ?

OpenStudy (phi):

no. it's clear. can you pick out the answer?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

(2^6)^-5??

OpenStudy (phi):

yes. the answer is 2^(-30) the way to get that the long way is "flip" and make -5 +5 \[ \left(2^6\right)^{-5}= \frac{1}{\left(2^6\right)^{5}}= \frac{1}{2^6\cdot 2^6 \cdot 2^6\cdot 2^6\cdot 2^6}\]

OpenStudy (anonymous):

ok

OpenStudy (phi):

2^6 is 2*2*2*2*2*2 and there are 5 of those, which adds up to 2 times itself 30 times.

OpenStudy (anonymous):

so when its multiplied up its 2^-(30) because 6x5 is 30

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok ya

OpenStudy (phi):

if you learn to do it the "hard way" it should make sense. and the short cuts will be easier to understand.

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

alrighty thanks for the awesome help again i got 2 left